Circular No. 8581

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005dg IN ESO 420-3 L. A. G. Monard, Pretoria, S. Africa, reports his discovery of an ap-

parent supernova (mag ∼ 16.6 ± 0.2) on three unfiltered CCD images taken on 2005 Aug. 5.14 UT with a 0.30-m Schmidt-Cassegrain reflector; a confirming image was obtained on Aug. 6.136 (the object’s magnitude then ∼ 16.5). SN 2005dg is located at α = 4^{h}07^{m}46^{s}.25, δ = −29^{o}51′28′′.1 (equinox 2000.0), which is 7′′ east and 6′′ north of the nucleus of ESO 420-3. Nothing is visible at this location on the Digitized Sky Survey (limiting red mag 20.5) or on an image taken on 2005 July 20.130 by Monard (limiting red mag 18.2).

SUPERNOVA 2005df IN NGC 1559 M. Salvo and B. Schmidt, Australian National University (ANU); and

M. Owers, University of New South Wales, report that preliminary reduction of a spectrum (range 390–700 nm), obtained on Aug. 5.83 UT with the ANU 2.3-m telescope (+ Double-Beam Spectrograph), shows SN 2005df (IAUC 8580) to be a somewhat-unusual young type-Ia supernova. Crosscorrelation with a library of template spectra suggests a similarity between this object and SN 1986G when a few days before maximum light (Phillips et al. 1987, PASP 99, 592). Adopting the NED recession velocity for the host galaxy (from Koribalski et al. 2004, A.J. 128, 16) the expansion velocity of the Si-II 635.5-nm line is around 15300 km/s.

SUPERNOVA 2005de IN UGC 11097 R. J. Foley, D. Perley, J. S. Bloom, University of California, Berkeley;

and J. X. Prochaska, University of California, Santa Cruz, report that inspection of CCD spectra (range 390–1000 nm), obtained on Aug. 4.34 UT with the Keck II 10-m telescope (+ ESI), shows that SN 2005de (cf. IAUC 8580) is of type Ia, resembling SN 1992A (Kirshner et al. 1993, Ap.J. 415, 589) at a few days before maximum brightness. Adopting a redshift of 4481 km/s (from Na D absorption lines in the spectrum), the expansion velocity of Si II (rest 635.5 nm) is ∼ 13300 km/s.

V1188 SCORPII Additional visual magnitude estimates (cf. IAUC 8575): July 30.93

UT, 9.6 (J. G. de S. Aguiar, Campinas, Brazil); Aug. 3.90, 10.6 (J. Carvajal, Avila, Spain).

2005 August 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green